# Exact WKB in all sectors I: Potentials with degenerate saddles

### v1

@@ 4.1 Transition between different sectors
Let us start with the Airy-type EWKB. The Stokes geometry remains the same for $u\in\left(0,u_{\mathrm{T}}\right)$. Without an analytic continuation, as $u$ passes through $u=u_{\mathrm{T}}$, the simple turning points first merge and then scatter off. As a result, we end up with a different Stokes geometry for the sector $u>u_{\mathrm{T}}$. See Fig. 10.
@@
We first observe that for the sector $u<u_{\mathrm{T}}$, the diagrams for $\theta_{\delta}=0^{\mp}$ are equivalent to the ones with the analytic continuation of $g$ with $\mathrm{Im}\,g=0^{\pm}$. (See e.g. Sueishi:2021xti .) As a result the exact quantization conditions for $\theta_{\delta}=0^{\mp}$ is expected to recover the known results in Sueishi:2021xti . In addition to that in Fig. 10, it is clear that with the analytic continuations $\theta_{\delta}:0^{\mp}\rightarrow\pi^{\pm}$ we introduced, the paths connecting $x=-\frac{\pi}{2}$ to $x=\frac{\pi}{2}$ pass through the same type of Stokes lines and branch points in the same order. Therefore, the transition matrices in both sectors are the same:
@@
$D_{\theta_{d}=0^{\mp}}=D_{\theta_{d}=\pm\pi^{\pm}}=2\cos\theta-\sqrt{\Pi_{A}^{-1}\Pi_{B}^{-1}}-\sqrt{\Pi_{A}\Pi_{B}^{-1}}-\sqrt{\Pi_{A}^{\mp 1}\Pi_{B}}=0.$ (128)
For $u<u_{\mathrm{T}}$, it is known that this is equal to the Zinn-Justin and Jentschura exact quantization condition Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw which becomes apparent when the Weber-type approach is incorporated. For completeness we write it using the explicit expressions in (117) and (118):
@@
In addition to that, in Basar:2015xna , it was argued that the transition between the different sectors of this real spectrum is smooth. This is the case after the Borel summation and the cancellation of the ambiguities. In the EWKB framework, this procedure is equivalent to taking the median summation of the exact quantization condition (128), which we denote by $D^{\mathrm{med}}$. Then, in view of the discussions in Basar:2015xna , we expect $D^{\mathrm{med}}$ stays intact upon transition between the two sectors which we discuss next [^8] We note that in this paper, we don’t introduce the median summation in connection with the directional Borel summation procedure. We refer reader to Kamata:2023opn for a detailed introduction and pedagogical discussion..
@@
The main object we use in the following discussion is the Stokes automorphism operation, $\mathfrak{S}$, which encodes the discontinuity when a Stokes diagram is degenerate. In that sense, $\mathfrak{S}$ relates the two exact quantization conditions connected by an analytic continuation which breaks the degeneracy. For example, in our discussion, these are the cases $\theta=0^{\pm}$ for $u<u_{\mathrm{T}}$ sector. The median summation, on the other hand, is equivalent to acting half-Stokes automorphism $\mathfrak{S}^{\pm 1/2}$ on the exact quantization conditions. Since the exact quantization condition contains both perturbative and non-perturbative ambiguities, this procedure is equivalent to the cancellation of the ambiguities.
Below the Barrier Top: Let us first focus on the $u<u_{\mathrm{T}}$ sector. For the diagrams in Fig. 9(a), this relationship is established via to the discontinuity of the $A$-cycle. At $\theta_{\delta}=0$, it is expressed as
@@
The median summation in this perspective can be obtained by $\mathfrak{S}^{\pm 1/2}D_{\theta_{\delta}=0^{\mp}}$. From (131), it is clear that the same median summation can be obtained from both $D_{\theta_{\delta}=0^{\mp}}$. Then, using (128) and (130), we get
@@
The resulting median quantization condition $D^{\mathrm{med}}_{u<\frac{1}{2}}$ is manifestly real. To see this reality, let us denote $\mathcal{C}$ as complex conjugation operator. Then, since $\mathcal{C}\left[\Pi_{A}\right]=\Pi^{-1}_{A}$ and $\mathcal{C}\left[\Pi_{B}\right]=\Pi_{B}$, it is clear that
@@
The reality of $D^{\mathrm{med}}_{u<u_{\mathrm{T}}}$ means that the trans-series of $u$ is also real and ambiguity free DDP2 . As we discussed above, this is one of the crucial aspects of the resurgence analysis of the periodic potential Zinn-Justin:1981qzi ; Zinn-Justin:1983yiy ; Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw ; Dunne:2013ada ; Dunne:2014bca . The EWKB analysis makes it manifest to all order. We can now extend it to the sector $u>u_{\mathrm{T}}$ using the analytic continuation we introduced.
@@
This is a result of the analytic continuation $u\rightarrow u\pm i\varepsilon$ and it is equivalent to $\theta_{\delta}:0^{\pm}\rightarrow\pm\pi^{\mp}$ in our discussion. Since we search for a smooth transition between the sectors $u<u_{\mathrm{T}}$ and $u>u_{\mathrm{T}}$, we need to adjust the cycles $A$ and $B$ accordingly. The analytic continuation in (136) suggests that for $\theta_{\delta}:0^{\pm}\rightarrow\pm\pi^{\mp}$, these transitions are incorporated into the exact quantization conditions via
@@
Now we can perform the median summation at $\theta_{\delta}=\pm\pi$: For convenience in the computations, we first invert the $B$-cycle and rewrite $D_{\theta_{\delta}=\pm\pi^{\mp}}$ in (128) as
@@
Then, the equality of (134) and (141) shows that the median summation of the exact quantization conditions in the sectors $u<u_{\mathrm{T}}$ and $u>u_{\mathrm{T}}$ are linked to each other by continuous analytic continuations and appropriate transformations. Note that for $u>u_{\mathrm{T}}$, $D^{\mathrm{med}}_{u>u_{\mathrm{T}}}$ is manifestly real as it is for $u<u_{\mathrm{T}}$. This presents the smooth connection of the resurgent trans-series of $u$ throughout the spectrum. For first few non-perturbative orders in the trans-series, this was previously discussed in Basar:2015xna using all order WKB expansion. With EWKB method, we make this conjecture manifestly exact.
@@ 4.2 The duality
$S^{u\sim 1}_{\mathrm{I}}=\int_{0}^{\pi}{\mathrm{d}}x\,\sqrt{2\left(V(x)-u_{\mathrm{T}}\right)}=2\,i\,,$ (147)
@@ 5 Symmetric double-well potential
for both sectors, we perform an exact quantization in a parallel manner to the periodic potential. Note that we introduce the simple rescaling by $\frac{1}{16}$ to keep the behaviour around a saddle point $x_{0}$ as $\frac{1}{2}\left(x-x_{0}\right)$. This doesn’t alter arguments on the analytic continuation of the classical action in Section 3 which we continue using as guidelines.
@@
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{g}a^{\mathrm{D}}(u,g)}=e^{-\mathcal{G}(\frac{u}{g},g)}\frac{\sqrt{2\pi}\left(\frac{g}{\mathcal{C}_{\mathrm{DW}}}\right)^{-\mathcal{F}\left(\frac{u}{g},g\right)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}\left(\frac{u}{g},g\right)\right)}\,,$ (154)
@@
    $\Pi_{\tilde{A}}$ $=e^{-\frac{2\pi i}{\hbar}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{1-u}{g},ig)}\frac{\sqrt{2\pi}\left(\frac{g}{\mathcal{C}_{4}}\right)^{-\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)}\,,$ (155)
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{\hbar}a^{\mathrm{D}}(u,g)}=e^{-2\pi i\,\left(\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)}\,,$ (156)
@@
and it represents an unstable quartic oscillator. Note that around $x=0$, $V_{\mathrm{D}}\simeq\frac{x^{2}}{4}+O(x^{3})$ which is the case for $\omega=\frac{1}{\sqrt{2}}$ in the Weber-type construction we discussed in Section 2.2.
Like periodic potential, when $u<u_{\mathrm{T}}$ , the symmetric double-well and its resurgence structure is well-studied via several methods Zinn-Justin:2004vcw ; Dunne:2013ada ; Dunne:2014bca ; Behtash:2015loa as well as the EWKB method Sueishi:2020rug . The P-NP relation is also well-known in this sector Alvarez1 ; Alvarez2 ; Alvarez3 ; Dunne:2013ada ; Dunne:2014bca . For our particular choice in (151), it is written as
$\mathcal{G}(\tilde{u},g)=-\int\frac{{\mathrm{d}}g}{g}\left[\frac{S_{\mathrm{I}}}{g}\frac{\partial\tilde{u}\left(\mathcal{F},g\right)}{\partial\mathcal{F}}+\mathcal{F}\right]\,,$ (159)
@@
The $u>u_{\mathrm{T}}$ sector, on the other hand, is less studied. In DDP2 , the expressions in (155) and (156) were briefly discussed but its explicit connection to the $u>u_{\mathrm{T}}$ sector was left out.
Finally, note that contrary to the case for the periodic potential, the dual theory of the double-well potential is not the same as itself. Therefore, it is expected that the P-NP of the dual theory would be different than P-NP relation in (159). In this particular case, a similar discussion was carried out in Codesido:2017dns where the duality transformation was performed for the free energy (prepotential) of the quantum system and the P-NP relation for the quartic oscillator was indicated in the dual theory. In our discussion, on the other hand, we perform a transformation for $\mathcal{F}$ and $\mathcal{G}$ which leads to the dual P-NP relation to (159) directly.
@@ 5.1 Transition between different sectors
Then, $\theta_{\delta}=0^{\pm}$ are the two analytic continuations which break the degeneracy in $u<u_{\mathrm{T}}$ sector and the analytic continuations $\theta_{\delta}:0^{\pm}\rightarrow\pm\pi^{\mp}$ connect the Stokes geometries of the two sectors continuously. Note that in this case, only two Stokes diagrams emerging from simple turning points are subject to the analytic continuation as crossing the level $u=u_{\mathrm{T}}$. Two possible analytic continuations of these diagrams are illustrated in Fig. 12.
@@
Figure 12: Stokes diagrams of the symmetric double-well potential. (a) The diagrams in $u<u_{\mathrm{T}}$ sector. The analytic continuations $\theta_{\delta}=0^{\mp}$ are equivalent to $\arg g=0^{\pm}$. (b) The diagrams in $u>u_{\mathrm{T}}$ sector. The diagrams at phases $\theta_{\delta}=\mp\pi^{\pm}$ are continuously connected to $\theta_{\delta}=0^{\pm}$.
Despite the significant change in the Stokes geometry, Fig. 12 shows that the transition matrices for $\theta_{\delta}=0^{\mp}$ are equal to $\theta_{\delta}=\pm\pi^{\mp}$ since the path connecting the infinities passes through the Stokes lines and branch cuts in the same order. The corresponding transition matrices are given as
@@
$T_{\theta_{\delta}=0^{+}}$ $=M_{\mathrm{A}}^{+}M_{\mathrm{A}}^{-}M_{\mathrm{A}}^{\mathrm{B}}N_{\mathrm{A}}^{3,4}M_{\mathrm{A}}^{+}N_{\mathrm{A}}^{2,3}M_{\mathrm{A}}^{+}M_{\mathrm{A}}^{-}M_{\mathrm{A}}^{\mathrm{B}}N_{\mathrm{A}}^{1,2}M_{\mathrm{A}}^{+}=T_{\theta_{\delta}=\pi_{-}}\,.$ (164)
@@
Note that the square root prefactors don’t play any role in physical interpretation of the exact quantization procedure and to obtain the quantized spectrum, the expressions in the brackets are set to zero. However, we observe $\sqrt{\Pi_{A}^{-2}\Pi_{B}^{-1}}$ factors are important in order to capture the precise relation between $D_{\theta_{\delta}=0^{\pm}}$ via a Stokes automorphism and in the computation of the median summations at $\theta_{\delta}=0$ as well as at $\theta_{\delta}=\pm\pi$. Therefore, we choose to leave the prefactor in the quantization conditions.
For $u<u_{\mathrm{T}}$ sector, (165) and (166) equal to ZJJ exact quantization conditions Zinn-Justin:2004vcw ; Dunne:2013ada ; Dunne:2014bca ; Sueishi:2020rug which becomes apparent when (153) and (154) are used to write them as
@@
As we mentioned above, the geometry of the $u>u_{\mathrm{T}}$ sector is the same as the stable quartic oscillator Bucciotti:2023trp ; Kamata:2023opn ; Kamata:2024tyb . Indeed, it is possible to interpret (165) and (166) as the exact quantization conditions for the quartic oscillator [^10] There are differences between (165) and (166) and the quantization conditions in Bucciotti:2023trp ; Kamata:2023opn ; Kamata:2024tyb . This is simply due to the branch cut choices in the Stokes diagrams which leads to different interpretations of $\Pi_{A}$ and $\Pi_{B}$ in different settings. (See Appendix A.). The main difference with the quartic oscillator and $u>u_{\mathrm{T}}$ sector of the symmetric double-well, on the other hand, becomes apparent when the Weber-type approach is incorporated: The associated Stokes diagram in the double-well potential for $u=u_{\mathrm{T}}+g\tilde{u}$ is given in Fig.13 and this is quite different than the Stokes diagram for the quartic oscillator at $u=0+g\tilde{u}$. (See Fig. 6 in Kamata:2023opn .)
@@
Now we turn our attention to the median summation of the exact quantization conditions in (165) and (166), and the smooth transition of the spectrum through $u=u_{\mathrm{T}}$.
@@
Then, we get the median summation of the quantization conditions as
$D^{\mathrm{med}}_{u<1}$ $=\mathfrak{S}_{0}^{1/2}D_{\theta_{\delta}=0^{-}}=\mathfrak{S}_{0}^{-1/2}D_{\theta_{\delta}=0^{+}}$ (170)
@@
Now, we perform the median summation at $\theta_{\delta}=\pm\pi$. As in Section 4, we start with inverting $B$-cycle and rewrite
@@
where we invert $B^{-1}$ cycle back to $B$-cycle. Finally, using (173), we get the same median summation for both cases as
$D^{\mathrm{med}}_{u>1}=\sqrt{\Pi_{\tilde{A}}^{-2}\Pi_{B}^{-1}}\,\left[\left(1+\Pi_{\tilde{A}}^{2}\right)\sqrt{1+\Pi_{B}}+2\Pi_{\tilde{A}}\right]\,.$ (178)
Again, we emphasize the role of the square-root prefactors in the computations leading the same result for both analytic continuations although they don’t have any physical implications. For example, when imposing (173) in (177), they led to the cancellations of $\Pi^{\mp}_{B}$ terms which appear differently at $\theta_{\delta}=+\pi$ and $\theta_{\delta}=-\pi$. Therefore, we observe that there is a single median summation in the $u>u_{\mathrm{T}}$ region and it has exactly the same form as (171), the median summation in $u<u_{\mathrm{T}}$ sector.
In addition to the equality of the median summations (171) and (178), they are also manifestly real, which indicates the reality of the spectrum for all $u$. As a result, we conclude that there is a smooth continuous transition of the spectrum between the two sectors and the resurgence structure stays intact in entire spectrum.

### v2

@@ 4.1 Transition between different sectors
Let us start with the Airy-type EWKB. For $u\in\left(0,u_{\mathrm{T}}\right)$, The Stokes geometry remains the same. Without an analytic continuation, as $u$ passes through $u=u_{\mathrm{T}}$, the simple turning points first merge and then scatter off. As a result, we end up with a different Stokes geometry for the sector $u>u_{\mathrm{T}}$. (See Fig. 10).
@@
We first observe that for the sector $u<u_{\mathrm{T}}$, the diagrams for $\theta_{\delta}=0^{\mp}$ are equivalent to the ones with the analytic continuation of $g$ with $\mathrm{Im}\,g=0^{\pm}$. (See e.g. Sueishi:2021xti .) As a result the exact quantization conditions for $\theta_{\delta}=0^{\mp}$ is expected to recover the known results in Sueishi:2021xti . In addition to that in Fig. 10, it is clear that with the analytic continuations $\theta_{\delta}:0^{\mp}\rightarrow\mp\pi^{\pm}$ we introduced, the paths connecting $x=-\frac{\pi}{2}$ to $x=\frac{\pi}{2}$ pass through the same type of Stokes lines and branch points in the same order. Therefore, the transition matrices in both sectors are the same:
@@
$D_{\theta_{\delta}=0^{\mp}}=D_{\theta_{\delta}=\pm\pi^{\pm}}=2\cos\theta-\sqrt{\Pi_{A}^{-1}\Pi_{B}^{-1}}-\sqrt{\Pi_{A}\Pi_{B}^{-1}}-\sqrt{\Pi_{A}^{\mp 1}\Pi_{B}}=0.$ (128)
For $u<u_{\mathrm{T}}$, it is known that this is equal to the ZJJ’s exact quantization condition Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw which becomes apparent when the Weber-type approach is incorporated. For completeness we write it using the explicit expressions in (117) and (118):
@@
In addition to that, in Basar:2015xna , it was argued that the transition between the different sectors of this real spectrum is smooth. This is the case after the Borel summation and the cancellation of the ambiguities. In the EWKB framework, this procedure is equivalent to taking the median summation of the exact quantization condition (128). In this paper, we don’t introduce the median summation in connection with the directional Borel summation procedure. We summarize basic facts in the following and refer reader to Kamata:2023opn for a detailed introduction and pedagogical discussion:
The main object we use in our discussion is the Stokes automorphism operation, $\mathfrak{S}$, which encodes the discontinuity when a Stokes diagram is degenerate. In that sense, $\mathfrak{S}$ relates two exact quantization conditions connected by an analytic continuation which breaks the degeneracy. For example, for $u<u_{\mathrm{T}}$ sector, around $\theta_{\delta}=0$, the exact quantization conditions $D_{\theta_{\delta}=0^{\pm}}$ are related via $\mathfrak{S}$, which we elaborate later.
The median summation, on the other hand, is equal to performing the Borel summation after removing the discontinuity from the exact quantization condition via the action of half-Stokes automorphism $\mathfrak{S}^{\pm 1/2}$. At $\theta_{\delta}=0$, the resulting object is expressed by
$\mathcal{S}_{\pm}\left[\mathfrak{S}^{\mp 1/2}D_{\theta_{\delta}=0^{\pm}}\right]\,,$
where $\mathcal{S}_{\pm}$ is the directional Borel summation. This procedure leads to an ambiguity free and real spectrum and it is equivalent to the cancellation of the ambiguities in a more direct but formal manner.
Note that although, it is desirable to obtain the Borel summed expressions for the spectrum, it is almost never a tractable process. In the EWKB framework, on the other hand, performing $\mathfrak{S}^{\pm 1/2}$ on the exact quantization conditions brings them to manifestly real forms [^10] For explicit examples on the disappearance of the imaginary terms in the trans-series arising from the exact quantization conditions, we refer to Kamata:2021jrs . For simple toy model examples, see also Sections III.A.1 and III.A.2 in Kamata:2023opn . which indicates the reality of the spectrum DDP2 . In our discussion, we define the exact quantization condition at a phase $\theta_{\delta}$ with discontinuities removed as $D^{\mathrm{med}}_{\theta_{\delta}}$ and call it “median QC”. In view of the discussions in Basar:2015xna , we expect $D^{\mathrm{med}}$ stays intact upon transition between the two sectors.
@@
Now, we discuss how $D^{\mathrm{med}}$ stays intact in both sectors of the periodic potential or equivalently at phases $\theta=0$ and $\theta=\pm\pi$. Note that along with the Stokes automorphism, the preservation of the quantization condition is provided by an appropriate analytic continuation of the $A$-cycle upon transition barrier in a compatible way with the analytic properties of the corresponding Airy-type WKB action.
Below the Barrier Top: Let us first focus on the $u<u_{\mathrm{T}}$ sector. For the diagrams in Fig. 9(a), this relationship is established via to the discontinuity of $\Pi_{A}$. At $\theta_{\delta}=0$, it is expressed as
@@
The median QCs in this perspective can be obtained by $\mathfrak{S}^{\pm 1/2}D_{\theta_{\delta}=0^{\mp}}$. From (131), it is clear that the median QCs are the same for both $D_{\theta_{\delta}=0^{\mp}}$ and using (128) and (130), we get
@@
Note that $D^{\mathrm{med}}_{u<u_{\mathrm{T}}}$ is manifestly real. To see this, let us denote $\mathcal{C}$ as complex conjugation operator. Then, since $\mathcal{C}\left[\Pi_{A}\right]=\Pi^{-1}_{A}$ and $\mathcal{C}\left[\Pi_{B}\right]=\Pi_{B}$, it is clear that
@@
The reality of $D^{\mathrm{med}}_{u<u_{\mathrm{T}}}$ means that the trans-series of $u$ is also real and ambiguity free DDP2 . As we discussed above, this is one of the crucial aspects of the resurgence analysis of the periodic potential Zinn-Justin:1981qzi ; Zinn-Justin:1983yiy ; Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw ; Dunne:2013ada ; Dunne:2014bca . The EWKB analysis makes it manifest to all order. We can now extend it to the sector $u>u_{\mathrm{T}}$.
@@
Note that the analytic continuation of $a_{0}(u)$ is provided by $u\rightarrow u\pm i\varepsilon$ and it is equivalent to $\theta_{\delta}:0^{\pm}\rightarrow\pm\pi^{\mp}$ in our discussion. Then, since we search for a smooth transition between the sectors $u<u_{\mathrm{T}}$ and $u>u_{\mathrm{T}}$, we need to adjust the cycles $A$ and $B$ accordingly. The analytic continuation in (136) suggests that for $\theta_{\delta}:0^{\pm}\rightarrow\pm\pi^{\mp}$, these transitions are incorporated into the exact quantization conditions via
@@
Now we can compute the median QCs at $\theta_{\delta}=\pm\pi$: For convenience in the computations, we first invert the $B$-cycle and rewrite $D_{\theta_{\delta}=\pm\pi^{\mp}}$ in (128) as
@@
Then, the equality of (134) and (141) shows that the median QCs, so that median summations, of the exact quantization conditions in the sectors $u<u_{\mathrm{T}}$ and $u>u_{\mathrm{T}}$ are linked to each other by continuous analytic continuations and appropriate transformations. Note that for $u>u_{\mathrm{T}}$, $D^{\mathrm{med}}_{u>u_{\mathrm{T}}}$ is manifestly real as it is for $u<u_{\mathrm{T}}$. This presents the smooth connection of the resurgent trans-series of $u$ throughout the spectrum. For first few non-perturbative orders in the trans-series, this was previously discussed in Basar:2015xna using all order WKB expansion. With the EWKB method, we make this conjecture manifestly exact.
@@ 4.2 The duality
$S^{u\sim u_{\mathrm{T}}}_{\mathrm{I}}=\int_{0}^{\pi}{\mathrm{d}}x\,\sqrt{2\left(V(x)-u_{\mathrm{T}}\right)}=2\,i\,,$ (147)
@@ 5 Symmetric double-well potential
for both sectors, we perform an exact quantization in a parallel manner to the periodic potential. Note that we introduce the simple rescaling by $\frac{1}{16}$ to keep the behaviour around a saddle point $x_{0}$ as ±$\pm\frac{1}{2}\left(x-x_{0}\right)$. This doesn’t alter arguments on the analytic continuation of the classical action in Section 3 which we continue using as guidelines.
@@
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{g}a^{\mathrm{D}}(u,g)}=e^{-\mathcal{G}(\frac{u}{g},g)}\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{DW}}}\right)^{-2\mathcal{F}\left(\frac{u}{g},g\right)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}\left(\frac{u}{g},g\right)\right)\right]^{2}}\,,$ (154)
@@
    $\Pi_{\tilde{A}}$ $=e^{-\frac{2\pi i}{g}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{1-u}{g},ig)}\frac{\sqrt{2\pi}\left(\frac{g}{\mathcal{C}_{4}}\right)^{-\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)}\,,$ (156)
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{g}a^{\mathrm{D}}(u,g)}=e^{-2\pi i\,\left(\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)}\,,$ (157)
@@
and it represents an unstable quartic oscillator.
Like periodic potential, when $u<u_{\mathrm{T}}$ , the symmetric double-well and its resurgence structure is well-studied via several methods Zinn-Justin:2004vcw ; Dunne:2013ada ; Dunne:2014bca ; Behtash:2015loa as well as the EWKB method DDP2 ; DP1 ; Sueishi:2020rug . The P-NP relation is also well-known in this sector Alvarez3 ; Dunne:2013ada ; Dunne:2014bca . For our particular choice in (151), it is written as
$\mathcal{G}(\tilde{u},g)=-2\int\frac{{\mathrm{d}}g}{g}\left[\frac{S_{\mathrm{I}}}{g}\frac{\partial\tilde{u}\left(\mathcal{F},g\right)}{\partial\mathcal{F}}+\mathcal{F}\right]\,,$ (160)
@@
The $u>u_{\mathrm{T}}$ sector, on the other hand, was studied in DDP2 ; DP1 , where the computations for $\mathcal{F}^{\mathrm{D}}$, $\mathcal{G}^{\mathrm{D}}$ and $\mathcal{C}_{4}$ was performed to get the expressions in (156). However, the explicit relationship of (156) with the unstable quartic oscillator wasn’t discussed. This also left out the possibility of observing the $S$-duality in the EWKB framework.
Finally, note that contrary to the case for the periodic potential, the dual theory of the double-well potential is not the same. Therefore, it is expected that the P-NP of the dual theory would be different than the P-NP relation in (160). In this particular case, a similar discussion was carried out in Codesido:2017dns where the duality transformation was performed for the free energy (prepotential) of the quantum system and the P-NP relation for the quartic oscillator was indicated in the dual theory. In our discussion, on the other hand, we perform a transformation for $\mathcal{F}$ and $\mathcal{G}$ which leads to the dual P-NP relation to (160) directly.
@@ 5.1 Transition between different sectors
Then, $\theta_{\delta}=0^{\pm}$ are the two analytic continuations which break the degeneracy in $u<u_{\mathrm{T}}$ sector and the analytic continuations $\theta_{\delta}:0^{\pm}\rightarrow\pm\pi^{\mp}$ connect the Stokes geometries of the two sectors continuously. Note that in this case, only two of the Stokes diagrams emerging from simple turning points are subject to the analytic continuation as crossing the level $u=u_{\mathrm{T}}$. Two possible analytic continuations of these diagrams are illustrated in Fig. 12.
@@
Figure 12: Stokes diagrams of the symmetric double-well potential. (a) The diagrams in $u<u_{\mathrm{T}}$ sector. The analytic continuations $\theta_{\delta}=0^{\mp}$ are equivalent to $\arg g=0^{\pm}$. (b) The diagrams in $u>u_{\mathrm{T}}$ sector. The diagrams at phases $\theta_{\delta}=\mp\pi^{\pm}$ are continuously connected to $\theta_{\delta}=0^{\mp}$.
Despite the significant change in the Stokes geometry, Fig. 12 shows that the transition matrices for $\theta_{\delta}=0^{\pm}$ are equal to $\theta_{\delta}=\pm\pi^{\mp}$ since the path connecting the infinities passes through the Stokes lines and branch cuts in the same order. The corresponding transition matrices are given as
@@
$T_{\theta_{\delta}=0^{+}}$ $=M_{\mathrm{A}}^{+}M_{\mathrm{A}}^{-}M_{\mathrm{A}}^{\mathrm{B}}N_{\mathrm{A}}^{3,4}M_{\mathrm{A}}^{+}N_{\mathrm{A}}^{2,3}M_{\mathrm{A}}^{+}M_{\mathrm{A}}^{-}M_{\mathrm{A}}^{\mathrm{B}}N_{\mathrm{A}}^{1,2}M_{\mathrm{A}}^{+}=T_{\theta_{\delta}=\pi^{-}}\,.$ (165)
@@
Note that the square root prefactors don’t play any role in the physical interpretation of the exact quantization procedure and to obtain the quantized spectrum, the expressions in the brackets are set to zero. However, we observe that $\sqrt{\Pi_{A}^{-2}\Pi_{B}^{-1}}$ factors are important in order to capture the precise relation between $D_{\theta_{\delta}=0^{\pm}}$ via a Stokes automorphism and in the computation of the median summations at $\theta_{\delta}=0$ as well as at $\theta_{\delta}=\pm\pi$. Therefore, we choose to leave the prefactor in the quantization conditions.
For $u<u_{\mathrm{T}}$ sector, (166) and (167) equal to ZJJ exact quantization conditions Zinn-Justin:2004vcw ; Dunne:2013ada ; Dunne:2014bca ; Sueishi:2020rug which becomes apparent when (153) and (LABEL:BcycleBelow_DW) are used to write them as
@@
As we mentioned above, the geometry of the $u>u_{\mathrm{T}}$ sector is the same as the stable quartic oscillator Bucciotti:2023trp ; Kamata:2023opn . Indeed, it is possible to interpret (166) and (167) as the exact quantization conditions for the quartic oscillator [^12] There are differences between (166) and (167) and the quantization conditions in Bucciotti:2023trp ; Kamata:2023opn . This is simply due to the branch cut choices in the Stokes diagrams which leads to different interpretations of $\Pi_{A}$ and $\Pi_{B}$ in different settings. (See Appendix A.). The main difference with the quartic oscillator and $u>u_{\mathrm{T}}$ sector of the symmetric double-well, on the other hand, becomes apparent when the Weber-type approach is incorporated: The associated Stokes diagram in the double-well potential for $u=u_{\mathrm{T}}+g\tilde{u}$ is given in Fig.13 and this is quite different than the Stokes diagram for the quartic oscillator at $u=0+g\tilde{u}$. (See Fig. 6 in Kamata:2023opn .)
@@
Now we turn our attention to the median summations of the exact quantization conditions in (166) and (167), and the smooth transition of the spectrum across $u=u_{\mathrm{T}}$. As in the periodic potential case, we present the preservation of the median QC across the barrier top provided by appropriate analytic continuations.
@@
Then, we get the median QC as
@@
Now, we compute the median QCs at $\theta_{\delta}=\pm\pi$: As in Section 4, we start with inverting $B$-cycle and rewrite
@@
where we invert $B^{-1}$ cycle back to $B$-cycle. Finally, using (174), for both cases, we get
$D^{\mathrm{med}}_{u>u_{\mathrm{T}}}=\sqrt{\Pi_{\tilde{A}}^{-2}\Pi_{B}^{-1}}\,\left[\left(1+\Pi_{\tilde{A}}^{2}\right)\sqrt{1+\Pi_{B}}+2\Pi_{\tilde{A}}\right]\,.$ (179)
Again, we emphasize the role of the square-root prefactors in the computations leading the same result for both analytic continuations although they don’t have any physical implications. For example, when imposing (174) in (178), they led to the cancellations of $\Pi^{\mp}_{B}$ terms which appear differently at $\theta_{\delta}=+\pi$ and $\theta_{\delta}=-\pi$. Therefore, we observe that there is a single median QC, in the $u>u_{\mathrm{T}}$ region and it has exactly the same form as (172), the median QC in $u<u_{\mathrm{T}}$ sector.
In addition to the equality of the median QCs (172) and (179), they are also manifestly real, which indicates the reality of the spectrum for all $u$. As a result, we conclude that there is a smooth continuous transition of the spectrum between the two sectors and the resurgence structure stays intact for the entire spectrum.